# AUDETER: A Large-scale Dataset for Deepfake Audio Detection in Open Worlds

### v1

@@ 3.3. Synthetic Audio Generation Process
Our synthetic audio is generated through two main processing pipelines: text-to-waveform synthesis via TTS systems and voice-to-voice conversion via vocoders, which constitute our two collections. For text-to-waveform generation, on the Common Voice, People’s Speech and the MLS subsets, we feed the transcripts to the TTS systems to generate corresponding synthetic speech. On the In-the-Wild dataset, we use OpenAI Whisper to generate transcripts of the bona-fide audio and then generate synthetic speech using the identical approach as above. For the vocoding collection, we apply the selected vocoders to the real audio from the subsets. It is worth noting that the synthetic patterns from the two collections differ significantly: TTS systems encode speaker references and semantic priors in the model pretraining, with many leveraging LLMs to generate semantic tokens. Vocoders encode their pretraining patterns as well as potential artefacts from the original audio.
@@ 3.4. Data Quality Assessment
To ensure the quality of our generated audio in terms of comprehensibility and naturalness indistinguishable from human speech, we throughly evaluate their intelligibility and naturalness.
@@ 3.4.1. Intelligibility Assessments
As a common quality measure for audio generation, intelligibility evaluation focuses on assessing whether the generated audio can preserve textual transcript information (Stevens et al., 2005). To make processing millions of audio files feasible, we adopt an automated evaluation approach, for which we employ a state-of-the-art Automatic Speech Recognition (ASR) model Whisper Large V3 (Radford et al., 2022) to convert generated audio clips to transcript and measure their semantic similarities with the ground truth transcript. We adopt four widely used evaluation metrics: WER Similarity (1-WER), Word Overlap, BLEU, and Exact Match, where WER focuses on transcription accuracy at the word level, Word Overlap measures lexical similarity, BLEU evaluates overall text quality, and Exact Match requires perfect transcription alignment. Figure 4 shows the WER similarity of the speech model on their generated audio for the People’s Speech Dataset. Due to space limitation, please refer to Section B for complete results for the other datasets. We found that the TTS models can produce better or at least comparable intelligibility to the original audio, and overall perform much better than the vocoders, demonstrating that the TTS models offer distinctly different and superior intelligibility patterns, showing the value of including them for comprehensive coverage.
@@ 3.4.2. Naturalness Assessment
The naturalness assessment focuses on human-like perceptual characteristics of generated audio. Mean Opinion Score (MOS) is a subjective quality metric where human listeners rate audio quality on a scale from 1 to 5 (higher is better). Due to the subjective nature of naturalness evaluation and our dataset’s large scale, we use the NISQA framework (Mittag and Möller, 2019) for automated perceptual quality prediction without requiring reference audio. The results of the average MOS score across all datasets are shown in Figure 5. Our findings here are similar to intelligibility, indicating that modern TTS systems substantially outperform vocoders in naturalness metrics, with several models approaching human-level perceptual quality.
3.5. Computational Cost
Given the scale of our dataset and varying efficiency of speech systems, we provide an estimate of our GPU usage. Dataset generation using NVIDIA A100 and H100 clusters consumed approximately 2000 GPU hours for synthetic audio and transcript generation for quality assessment.
@@ 4.1.1. Baseline Methods
We employ nine popular deepfake audio detection methods with publicly available implementations, including both crafted DNN models and detection models integrating pretrained backbones with scoring heads. Among the first category, we include RawNet2 (Tak et al., 2021a), RawGAT-ST (Tak et al., 2021b), AASIST(Jung et al., 2021), PC-Dart (Ge et al., 2021a), SAMO (Ding et al., 2023), Neural Vocoder Artifacts (NVA) (Sun et al., 2023) and Purdue M2 (Ren et al., 2025). For the second category, we adapt XLS-R + RawNet + Assist (XLS+R+A) (Tak et al., 2022) and XLS+SLS (Zhang et al., 2024).
@@ 4.1.2. Evaluation Metrics
Following (Tak et al., 2021a; Müller et al., 2022; Jung et al., 2021; Tak et al., 2022; Zhang et al., 2024), we employ the popular evaluation metric, Equal Error Rate (EER), to evaluate synthetic audio detection performance. EER represents the error rate where the false positive rate intersects the false negative rate, providing a balanced measure of detection accuracy that equally weighs errors in classifying both real and synthetic audio. Lower EER values indicate superior detection performance, with 0% representing perfect detection.
@@ 4.1.3. Implementation Details
For evaluating the pretrained detection models on AUDETER, their official code implementation are used with their best pretrained checkpoints for scoring. For data-centric supervised training experiments, we use the official implementations of the selected models and use a learning rate of 1e-06 and a batch size of 128 to train all models. Given the large number of speech system patterns in our dataset, we adopt a balanced batching strategy. At each iteration, we go through all training real samples in randomised order for half of the batch size and uniformly sample all fake training samples for the other half of the batch size. Due to space limitations, please refer to the appendix for detailed implementation details.
@@ 4.1.4. Evaluation Protocol
Since AUDETER contains multiple synthetic audio versions for each real audio subset, evaluation is performed iteratively. To evaluate a subset, real audio is combined with each synthetic version in turn, repeating for all synthetic patterns. For instance, a full evaluation round for the Common Voice subset involves combining its real audio with each TTS and vocoder version, resulting in 26 runs.
4.2. Results on AUDETER as Test Data
@@
We evaluate pretrained baseline models on AUDETER to examine their effectiveness for identifying novel deepfake patterns and diverse human speech corpora. The performance heatmap for all speech system are detailed in Figure 6. Please refer to Tables 17–24 in the appendix for the complete results on the In-the-Wild and MLS partition. We observe that the baseline methods struggle to achieve reasonable performance for most of the experiment.
@@
Vulnerability to Novel Deepfake Patterns. The baseline methods exhibit noticeable performance degradation when directly applied to audio samples from AUDETER, which are generated by recent speech synthesis models mostly not covered by existing datasets. We found no single detection model can consistently achieve usable performance (i.e., reasonably low EER) across all datasets, indicating that generalisation achieved by training on existing datasets is insufficient for open-world detection.
We observe the following interesting phenomena. First, more recent TTS systems create greater challenges for the detectors, suggesting that the development of speech synthesis systems, particularly after the widespread adoption of pretraining, generates audio with distinct acoustic characteristics. Second, domain shift in human speech significantly affects detection performance, as performance on synthetic audio from the same TTS system varies when the corresponding real audio comes from different corpora. Third, the two baselines leveraging pretrained backbones achieved relatively better overall performance. This could be attributed to their increased parameters or pretrained knowledge enabling better characterisation of real audio. In addition, we found that baseline methods tend to perform better on deepfake audio generated by vocoders. These results demonstrate that our dataset effectively challenges existing methods and highlight the importance of evaluating against a variety of domain shifts, consistent with our motivation.
@@
Effect of Speakers on Detectability. To analyse speaker effects on detection performance, we summarise baseline method EERs on synthetic audio from 5 OpenVoice versions across all datasets in Table 4. Significant variation across speakers indicates that vocal characteristics influence synthetic audio detectability beyond architectural differences.
4.3. Results on AUDETER as Training Data
AUDETER is a valuable resource for large-scale robust training. To demonstrate the effectiveness of leveraging diverse real and synthetic audio patterns as a data-centric approach for improving open-world detection performance, we perform experiments using two popular XLR-based and one DNN detection model for cross-domain evaluation. We also train the models using different combinations of AUDETER’s subsets to study the effect of training data composition. We adopt both In-the-Wild and AUDETER for open-world evaluation.
@@ 4.3. Results on AUDETER as Training Data
Model EER Purdue M2 (Ren et al., 2025) 79.75 PC-Dart (Ge et al., 2021a) 66.17 RawGAT-ST (Tak et al., 2021b) 52.60 AASIST (Jung et al., 2021) 43.02 RawNet2 (Tak et al., 2021a) 37.81 SAMO (Ding et al., 2023) 37.09 Wav2vec,HuBERT,Conformer & attention (Wang et al., 2023) 36.84 XLS-R & Res2Net (Yi et al., 2023) 36.62 MPE & SENet (Wang et al., 2024b) 29.62 NVA (Sun et al., 2023) 26.32 Spec & POI-Forensics (Pianese et al., 2022) 25.14 XLXS-R,WavLM,Hubert & Fusion (Yang et al., 2024) 24.27 XLR+R+A (Tak et al., 2022) 10.46 XLR-SLS (Zhang et al., 2024) 7.46 RawNet2 (Train on CV and PS subsets) 27.13 XLR+R+A (Train on CV and PS subsets) 5.05 XLR+SLS (Train on CV and PS subsets) 4.17
@@ 4.3.1. Cross Domain Generalisation from AUDETER to In-the-Wild
We choose ITW for evaluation because its synthesis methods and real audio sources are not explicitly discussed and differ from AUDETER, isolating against unintentional data leakage. We train XLS+R+A, XLR+SLS and RawNet2 using all audio from the Common Voice and People’s Speech subsets. Please note that the In-the-Wild subsets are excluded from any training combination to maintain the integrity of cross-domain evaluation.
@@
Table 5 compares the best performance of the XLR-A-R model, XLR-SLS and RawNet 2 model trained using all data from the Common Voice and People’s Speech Dataset) with the best performance of other detection models. For comparison, we use their reported results when available, otherwise we evaluate using their official weights. The models trained using AUDETER not only achieve lower EER compared to the baseline methods, but also demonstrate significant improvement compared to the performance using their offical pretrained weights. Specifically, the ERR are reduced from 7.46 to 4.17 for XLR+SLS, 10.46 to 5.05 for SLR+R+A, and 37.81 to 27.13 for Rawnet 2, achieving respectively 44.1%, 51.6%, and 28.2% reduction in EER. This demonstrates that AUDETER provides effective diversity of deepfake and real audio combinations for improving detection model training without requiring test domain knowledge through large-scale training. The performance gain comes from diverse patterns that implicitly improve generalisation rather than data augmentation techniques. Our synthesis models, selected for popularity and recency, avoid specifically mimicking any existing benchmark.
@@
ALL Unseen Model TTS VOC All TTS VOC All RawNet2 0.257 0.434 0.346 0.286 0.378 0.332 ASSIST 0.325 0.454 0.390 0.401 0.622 0.512 RawGAT-ST 0.266 0.475 0.371 0.346 0.447 0.397 PC-Dart 0.169 0.192 0.180 0.218 0.026 0.122 SAMO 0.492 0.506 0.499 0.471 0.475 0.473 NVA 0.492 0.551 0.522 0.450 0.408 0.429 Purdue-M2 0.438 0.564 0.501 0.537 0.715 0.626 XLR+R+A 0.365 0.192 0.279 0.346 0.239 0.292 XLR+SLS 0.299 0.148 0.224 0.275 0.164 0.220 XLR+R+A 0.035 0.018 0.026 0.028 0.015 0.021 XLR+SLS 0.016 0.013 0.014 0.010 0.026 0.018 RawNet 2 0.087 0.232 0.159 0.314 0.254 0.284
4.3.2. Cross Domain Generalisation Analysis within AUDETER
@@
To further demonstrate that large-scale training improves generalisation, we perform cross-domain evaluation using the same three detectors trained on Common Voice and People’s Speech subsets in two settings: (1) all TTS and vocoder versions, and (2) 5 selected TTS and 5 vocoders, then evaluate on the MLS subset. The first setting examines transfer learning under human speech domain shift only, while the second examines both domain shift and novel synthesis systems. We report average performance across all system versions for the first setting and across unseen systems for the second in Table 6. We observe that all three models trained using AUDETER achieve significant EER reduction compared to all pretrained models and their pretrained versions, demonstrating that large-scale training with diverse real and synthetic audio improves generalisation under domain shift. While all three methods achieve significant overall reduction, the XLR-based models achieve near-zero performance, much better than RawNet 2, highlighting the importance of having sufficient learning capacity to incorporate open-world knowledge. Training large models requires substantial data, which is what AUDETER provides.
Collection Subset Model Best EER TTS (All) CV PS XLR+R+A 11.39 XLR-SLS 11.59 Voc (All) CV PS XLR+R+A 6.18 XLR-SLS 6.09 TTS + Voc (All) CV XLR+R+A 23.71 XLR-SLS 12.13 5 TTS + 5 Voc CV PS XLR+R+A 5.52 XLR-SLS 5.18 TTS + Vocoder (All) CV + PS XLR+R+A 5.05 XLR-SLS 4.17
@@ 4.3.3. Effect of Synthetic Pattern Diversity in Large-scale Training
To show the benefit of including diverse audio patterns in large-scale training, we vary the subsets to train multiple models and compare their performance on the in-the-wild dataset, as described in Table 7. Specifically, we train the two architecture using: (1) both TTS and Vocoder collections with only the Common Voice subset, (2) only TTS collection with Common Voice and People Speech, and (3) only Vocoder collection with Common Voice and People Speech, using the same settings. For both models, it is not surprising that the models trained using both TTS and Vocoder Collections with samples from both Common Voice and Peoples Speech partitions yield the best performance, again demonstrating the benefits of training data diversity on open-world detection performance. On the contrary, we notice that some models trained with only the Common Voice subset, even using all TTS and vocoder fake audios, produce unusable performance, further highlighting the usefulness of having diverse real audio in training.
Although models trained using the Common Voice and Peoples Speech subsets from either the TTS or the vocoder partition achieve reasonable performance, this is less than when used in combination.
@@
To show that training with diverse data samples at greater scale can lead to better generalisation, we train the XLR+R+A model using different percentages of all data samples from the CV and PS subsets (10%, 20%, 40%, 60%, and 100%) and report the average performance of the top 3 best EER results on the In-the-Wild dataset. We observe consistent improvement as more data are used for training.
@@
A key consideration in developing open-world detection models is understanding the shared characteristics between different synthesis models and how training audio from one model generalises to others. The balanced and consistent structure of AUDETER facilitates this exploration.
@@
We train multiple XLR+R+A models using real audio from the Common Voice validation partition and synthetic audio generated by a single speech synthesis system with matching scripts. Specifically, we use deepfake samples from 5 TTS models (Fish-Speech, F5-TTS, SparkTTS, VITS, XTTS) and 5 vocoders (BigVGAN, UnivNet, MelGAN, HiFi-GAN, Vocos) and repeat the training process 10 times. Following the evaluation protocol in Sec. 4.1.4, we then evaluate the performance on different test sets. Due to space limitations, we report the average performance on the TTS and Vocoder Collection across three different test domains in Table 8. Please refer to Tables 25–36 for the complete results.
@@
Generalisation on Matched Synthetic Audio and Identical Human Speech. The results under the CV Val column show the average performance of models trained using single systems for each collection, where scripts and real audio are identical to training, evaluating the impact of synthetic patterns alone. Generalisation ability varies significantly across speech models. For instance, SparkTTS and Fish-Speech showing stronger generalisation among TTS systems. We observe that TTS systems and vocoders struggle to generalise to each other, with performance generally better within the same collection of models. Additionally, generalisation is more difficult for vocoders. These findings demonstrate the importance of including samples from both system types and diverse training systems for robust model training.
@@
Generalisation with Different Textual Content. We repeat the evaluation process on the Common Voice test partition, where textual information no longer matches but the same style of human voice is maintained. The results under the CV Test column shows marginally lower but comparable results compared to matched textual information, indicating that text content does not significantly affect generalisation.
@@
Generalisation across Domains and Text. We further evaluate using the People’s Speech test partition (i.e., the PS Test column in Table 8) to explore the effect of distribution shift with different text content. Performance significantly decreases compared to previous experiments, even for the same system used in training. This suggests that detection models are sensitive to real audio distribution shift, emphasising the importance of training data diversity.

### v2

@@ 3.3. Synthetic Audio Generation Process
AUDETER includes two synthetic audio generation pipelines: text to waveform synthesis using TTS systems and voice to voice conversion using vocoders. For TTS synthesis, waveform generation uses original transcripts when available, or transcripts obtained via Whisper (Radford et al., 2022) otherwise. For the vocoder collection, selected vocoders are applied directly to real audio samples. The two collections exhibit distinct synthetic patterns. TTS systems encode speaker references and semantic priors from model pretraining, often leveraging large language models, whereas vocoders reflect their pretraining characteristics and artefacts inherited from the original audio. AUDETER’s audio generation pipeline is illustrated in Fig. 8.
@@ 3.4. Data Quality Assessment
To ensure the quality of the generated audio in terms of intelligibility and naturalness, we conduct a thorough evaluation of its comprehensibility and perceptual quality.
@@ 3.4.1. Intelligibility Assessments
Intelligibility evaluates whether generated audio preserves the underlying transcript (Stevens et al., 2005). To enable large-scale evaluation, we adopt an automated approach using the Whisper Large V3 ASR model (Radford et al., 2022) to transcribe generated audio and compare it with the reference text. We report four standard metrics: WER similarity (1-WER), word overlap, BLEU, and exact match. Figure 3 (a) shows WER similarity results on the UC subset, with full results for other datasets reported in Section B. Overall, TTS models achieve comparable or better intelligibility than the original audio and consistently outperform vocoders, validating the quality of AUDETER’s synthetic samples.
@@ 3.4.2. Naturalness Assessment
Naturalness evaluates the perceptual similarity of generated audio to human speech. We use Mean Opinion Score (MOS), a standard subjective metric ranging from 1 to 5, and adopt the NISQA framework (Mittag and Möller, 2019) for automated MOS prediction due to the scale of the dataset. Figure 3 (b) reports average MOS scores across datasets. Similar to intelligibility, modern TTS systems substantially outperform vocoders in naturalness, with multiple models approaching human level perceptual quality.
@@ 4.1. Experimental Settings
4.1.1. Experiments Overview
Our experiments are organised into three main parts: evaluating pretrained detectors on AUDETER as a test benchmark (Sec. 4.2), training existing models on AUDETER for cross-domain evaluation on In-the-Wild and across AUDETER subsets (Sec. 4.3), and extending cross-domain analysis to DFADD and SONAR to examine learning under increasingly diverse deepfake sources, motivating and evaluating our proposed curriculum learning strategy (Sec. 4.4).
@@ 4.1.2. Baseline Methods
We employ nine popular deepfake audio detection methods with publicly available implementations, including both crafted DNN models and detection models that integrate pretrained backbones with scoring heads. Among the first category, we include RawNet2 (Tak et al., 2021b), RawGAT-ST (Tak et al., 2021a), AASIST(Jung et al., 2021), PC-Dart (Ge et al., 2021a), SAMO (Ding et al., 2023), Neural Vocoder Artifacts (NVA) (Sun et al., 2023) and Purdue M2 (Ren et al., 2025). For the second category, we adapt XLS-R + RawNet + Assist (XLS+R+A) (Tak et al., 2022) and XLS+SLS (Zhang et al., 2024).
@@ 4.1.3. Evaluation Metrics
4.1.4. Computational Cost
Given the scale of AUDETER and varying system efficiency, we estimate our GPU usage. Dataset generation on NVIDIA A100 and H100 clusters consumed approximately 2000 GPU hours for synthetic audio and transcript generation.
@@ 4.1.5. Implementation Details
Pretrained detectors are evaluated using their official implementations and released checkpoints. For supervised training on AUDETER, all models are trained with a learning rate of and batch size 128, using a balanced batching strategy with 64 real samples per epoch and 64 uniformly sampled fake samples per batch. All other settings follow the original setup. Importantly, all detection models operate solely on audio inputs and do not access textual transcripts or exploit script matching during either training or inference. For the proposed curriculum learning, we use the same settings for Stage 1 training. Due to space limitations, further details are provided in Sec.C.
@@ 4.1.6. Evaluation Protocol
Since AUDETER contains multiple synthetic versions per real subset, evaluation is performed by iteratively pairing real audio with one synthetic version at a time and averaging performance across runs, as detailed in Algorithm 1. For example, evaluating the CrowdSource subset yields 26 runs across all TTS and vocoder variants. The averaging can be applied per system, per collection, or across the full subset.
4.2. AUDETER as Test Data
@@
We evaluate pretrained baseline models on AUDETER to assess their ability to generalise to novel deepfake patterns and diverse human speech. Figure 4 summarises performance across speech systems, with complete results on the CE and AB subsets reported in Tables 19–26 in the appendix. Overall, baseline methods struggle to achieve robust performance across most settings.
@@
Performance on AUDETER. The baseline models exhibit substantial performance degradation under zero-shot detection on AUDETER, which contains audio generated by recent speech synthesis systems largely absent from existing datasets. No single model achieves consistently low EER across datasets, and the average EER of each model remains high, indicating that generalisation learned from existing benchmarks is insufficient for recent speech systems.
We observe several key trends: (i) detection accuracy drops notably on more recent TTS systems, suggesting increasing acoustic divergence with advances in speech synthesis; (ii) domain shifts in real speech further affect performance, as detectors trained on the same TTS system vary across different real speech corpora; and (iii) models with pretrained backbones achieve relatively stronger overall performance, likely due to improved modelling of real speech, while baseline methods generally perform better on vocoder generated audio. These observations confirm that AUDETER poses meaningful challenges to existing detectors and highlight the importance of evaluating robustness under diverse domain shifts.
@@
Effect of Speakers on Detectability. To analyse speaker effects on detection performance, we report baseline EERs on synthetic audio generated by five OpenVoice variants across all datasets in Table 18 in the appendix due to space limits. The observed variation across speakers indicates that vocal characteristics influence synthetic audio detectability beyond architectural differences.
4.3. AUDETER for Training Existing Models
To demonstrate AUDETER as a useful training resource for open-world detection, we conduct cross-domain evaluations using two popular XLR-based models and one DNN-based detector under their original architectures and training settings. AUDETER-trained models are evaluated on a standard benchmark (ITW) and across different subsets within AUDETER to assess how the training data influences generalisation under various domain shifts.
@@ 4.3. AUDETER for Training Existing Models
Model EER Purdue M2 (Ren et al., 2025) 79.75 PC-Dart (Ge et al., 2021a) 66.17 RawGAT-ST (Tak et al., 2021a) 52.60 AASIST (Jung et al., 2021) 43.02 RawNet2 (Tak et al., 2021b) 37.81 SAMO (Ding et al., 2023) 37.09 Wav2vec,HuBERT,Conformer & attention (Wang et al., 2023) 36.84 XLS-R & Res2Net (Yi et al., 2023) 36.62 MPE & SENet (Wang et al., 2024a) 29.62 NVA (Sun et al., 2023) 26.32 Spec & POI-Forensics (Pianese et al., 2022) 25.14 XLXS-R,WavLM,Hubert & Fusion (Yang et al., 2024) 24.27 XLR+R+A (Tak et al., 2022) 10.46 XLR-SLS (Zhang et al., 2024) 7.46 XLR+R+A (AUDETER (CS, UC), no aug) 5.05 XLR+SLS (AUDETER (CS, UC), no aug) 4.17 RawNet2 (AUDETER (CS, UC)) 27.13 XLR+R+A (AUDETER (CS, UC)) 2.03 XLR+SLS (AUDETER (CS, UC)) 1.87
@@ 4.3.1. Cross Domain Generalisation from AUDETER to In-the-Wild
ITW is the default benchmark for cross-domain DFAD and has no overlap with AUDETER in either real speech sources or synthesis methods. We train XLR+R+A, XLR+SLS, and RawNet2 using AUDETER’s CS and UC subsets. Table 4 compares the best performance of AUDETER-trained models with baseline DFAD methods. We use reported results when available and otherwise evaluate using the official released weights. AUDETER-trained achieve lower EER than baseline methods and substantially outperform their official pretrained versions. Specifically, EER is reduced from 7.46 to 1.87 for XLR+SLS, from 10.46 to 2.03 for XLR+R+A, and from 37.81 to 27.13 for RawNet2, corresponding to relative reductions of 74.9%, 80.6%, and 28.3%, respectively. Even without the default data augmentation used by existing methods, AUDETER-trained XLR-based models achieve EERs of 5.05 (XLR+SLS) and 4.17 (XLR+R+A), yielding relative reductions of 32.3% and 60.1% over their pretrained baselines. This shows that AUDETER improves detection training through diverse real and deepfake patterns, without relying on test-domain knowledge even without data augmentation.
@@
ALL Unseen Model TTS VOC All TTS VOC All RawNet2 0.257 0.434 0.346 0.286 0.378 0.332 AASIST 0.325 0.454 0.390 0.401 0.622 0.512 RawGAT-ST 0.266 0.475 0.371 0.346 0.447 0.397 PC-Dart 0.169 0.192 0.180 0.218 0.026 0.122 SAMO 0.492 0.506 0.499 0.471 0.475 0.473 NVA 0.492 0.551 0.522 0.450 0.408 0.429 Purdue-M2 0.438 0.564 0.501 0.537 0.715 0.626 XLR+R+A 0.365 0.192 0.279 0.346 0.239 0.292 XLR+SLS 0.299 0.148 0.224 0.275 0.164 0.220 XLR+R+A 0.035 0.018 0.026 0.028 0.015 0.021 XLR+SLS 0.016 0.013 0.014 0.010 0.026 0.018 RawNet 2 0.087 0.232 0.159 0.314 0.254 0.284
4.3.2. Cross-Domain Generalisation Analysis within AUDETER
@@
Cross-domain evaluation can be conducted within AUDETER by training on one or more subsets and testing on others. We evaluate generalisation on the AB subset using the three detectors trained on the CS and UC subsets under two settings: (i) training with all speech models’ variants, and (ii) training with variants from five selected TTS systems and five vocoders. The first setting isolates human speech domain shift, while the second additionally introduces unseen synthesis systems. Average performance for each setting is reported in Table 5. All three models trained on AUDETER achieve substantial EER reductions compared to their pretrained counterparts, demonstrating improved generalisation under domain shift, highlighting that AUDETER provides diverse and sufficiently rich training data for learning generalisable representations. Among them, XLR-based models achieve near zero EER and consistently outperform RawNet2, showing that sufficient learning capacity from large backbones is essential for encoding generalisable representations.
4.3.3. Robustness against Real-world Adversaries
Real-world factors such as background noise can act as adversarial perturbations that impact detection performance. We therefore evaluate the robustness of AUDETER under three widely observed noise conditions: white noise (), telephone effect (), and background noise (). Table 6 reports the performance differences () between evaluations with and without noise perturbations, where smaller values indicate stronger robustness. We observe that real-world adversaries generally increase EER across models, though the effect is mild. In comparison, models trained on AUDETER exhibit substantially smaller EER increases, demonstrating stronger robustness to these common perturbations.
Model XLR+R+A (ori.) 1.53 -1.10 14.82 5.08 XLR+SLS (ori.) 3.23 7.33 2.20 4.25 XLR+R+A (AUDE. CS+UC) 0.26 2.09 0.17 0.84 XLR+SLS (AUDE. CS+UC) 0.28 3.77 0.29 1.45
Collection Subset Model Best EER TTS (All) CS+UC XLR+R+A 11.39 XLR-SLS 11.59 Voc (All) CS+UC XLR+R+A 6.18 XLR-SLS 6.09 TTS + Voc (All) CS XLR+R+A 23.71 XLR-SLS 12.13 5 TTS + 5 Voc CS+UC XLR+R+A 5.52 XLR-SLS 5.18 TTS + Vocoder (All) CS+UC XLR+R+A 5.05 XLR-SLS 4.17
@@ 4.3.4. Effect of Synthetic Pattern Diversity in Large-scale Training
To show the benefit of including diverse audio patterns in large-scale training, we vary the subsets to train multiple models and compare their performance on the ITW dataset, as described in Table 7. Specifically, we train the two architecture using: (1) both TTS and Vocoder collections with only the CrowdSource subset, (2) only TTS collection with CrowdSource and US Congress, and (3) only Vocoder collection with CrowdSource and US Congress, using the same settings. For both models, it is not surprising that the models trained using both TTS and Vocoder Collections with samples from both CrowdSource and US Congress subsets yield the best performance, again demonstrating the benefits of training data diversity on open-world detection performance. In contrast, we notice that some models trained with only the CrowdSource subset, even using all TTS and vocoder fake audios, produce unusable performance, further highlighting the usefulness of having diverse real audio in training. Although models trained using the CrowdSource and US Congress subsets from either the TTS or the vocoder partition achieve reasonable performance, this is less than when used in combination.
@@
To show that training with diverse data samples at greater scale can lead to better generalisation, we train the XLR+R+A model using different percentages of all data samples from the CS and UC subsets (10%, 20%, 40%, 60%, and 100%) and report the average performance of the top 3 best EER results on the In-the-Wild dataset. We observe consistent improvement as more data are used for training.
@@
Understanding shared characteristics across synthesis models and their impact on cross-system generalisation is important for developing open-world DFAD models. The balanced and consistent structure of AUDETER facilitates this exploration. We train multiple XLR+R+A models using real audio from the CS subset and synthetic audio generated by a single speech synthesis system with matched scripts. Specifically, we consider five TTS models (Fish-Speech, F5-TTS, SparkTTS, VITS, XTTS) and five vocoders (BigVGAN, UnivNet, MelGAN, HiFi-GAN, Vocos), repeating the training process ten times for each system. Evaluation follows the protocol in Sec.4.1.6. Due to space constraints, Table 8 reports average performance across the TTS and Vocoder collections over three test domains, with full results provided in Tables 27–38.
@@
Generalisation on Matched Synthetic Audio and Identical Human Speech. The results in the CS Val column reflect average performance of models trained on single systems within each collection, where scripts and real audio match the training data, isolating the effect of synthetic patterns. Generalisation varies markedly across speech models, with SparkTTS and Fish-Speech exhibiting stronger transfer among TTS systems. Cross-collection generalisation between TTS systems and vocoders remains limited, with models performing better within the same collection, and vocoders generally exhibiting weaker generalisation. These results highlight the importance of including both system types and diverse synthesis models for robust detector training.
@@
Generalisation with Different Textual Content. We repeat the evaluation on the CS test partition, where textual content differs from training while the same speaker style is preserved. Results in the CS Test column are slightly lower but comparable to those with matched text, indicating that textual content has limited impact on generalisation.
@@
Generalisation across Domains and Text. We further evaluate on the UC test partition to assess the effect of distribution shift in real speech. Performance drops substantially compared to previous settings, even when using the same synthesis system as in training, indicating high sensitivity to real-audio domain shift and underscoring the importance of diverse real speech in training data.